# FB-RAG: Improving RAG with Forward and Backward Lookup

### v2

@@ Abstract
Traditional Retrieval-Augmented Generation (RAG) struggles with complex queries that lack strong signals to retrieve the most relevant context, forcing a trade-off between choosing a small context that misses key information and a large context that confuses the LLM. To address this, we propose Forward-Backward RAG (FB-RAG), a new training-free framework based on a simple yet powerful forward-looking strategy. FB-RAG employs a light-weight LLM to peek into potential future generations, using evidence from multiple sampled outputs to precisely identify the most relevant context for a final, more powerful generator. This improves performance without complex finetuning or Reinforcement Learning common in prior work. Across datasets from LongBench and Bench, FB-RAG consistently delivers strong results. Further, the performance gains can be achieved with reduced latency due to a shorter, more focused prompt for the powerful generator. On EN.QA dataset, FB-RAG matches the leading baseline with over % latency reduction or achieves an % performance improvement with a % latency reduction. Our analysis finds cases where even when the forward-looking LLM fails to generate correct answers, its attempts are sufficient to guide the final model to an accurate response, demonstrating how smaller LLMs can systematically improve the performance and efficiency of larger ones.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) shows immense promise in reducing hallucinations and improving generation performance fan2024survey; gao2023retrieval. RAG achieves strong results on diverse Question Answering (QA) tasks borgeaud2022improving; guu2020retrieval; asai2024self, general language tasks he2021fast; khandelwalgeneralization, and across numerous downstream applications liu2023multi; wu2024coral.
@@
In this work, we focus on the task of answering queries based on an already-provided large context. Traditional RAG efforts for this setup involve two steps zhao2024retrieval: 1) Retrieve important chunks by computing similarities with the query (based on a sparse or dense retriever and/or a reranker), 2) Feed the retrieved chunks along with the query to an LLM, which generates the answer. We refer to these approaches as backward-looking – looking back at the input query to score the context chunks. Such methods have been widely adopted in both academia and industry. However, standard methods struggle with complex queries that lack sufficient information to retrieve relevant chunks (see example in Figure 1). This challenge is difficult to manage in RAG, where retrieving too little risks missing key information and retrieving too much risks adding irrelevant content that can confuse the LLMs yu2024defense.
@@
To address this challenge, we design Forward-Backward RAG (FB-RAG) for studying the impact of an emerging yet underexplored idea – forward-looking or peeking into the LLM’s output generations to improve retrieval. FB-RAG generates the output in three stages: I) Recall-focused Retrieval, using an off-the-shelf retriever to extract a smaller, yet sufficiently large context, II) Precision-focused Retrieval, which either only relies on forward-looking by observing reasons and answers from a light-weight LLM to evaluate the context chunks (Ours-F) or relies on both forward and backward lookup (Ours-FB), and III) Generation, prompting a powerful LLM to get the final answer. Although prior work used related ideas to improve RAG with LLM-based feedback or confidence scores zhao2024longrag; sun2022recitation; wang2024speculative; yang2023prca; jiang-etal-2023-active, these methods typically propose complex fine-tuning or Reinforcement Learning strategies, and often assume access to external web search engines or rely on LLM’s own memory which is not suitable for many domain-specific practical settings. Instead, FB-RAG is a simple and effective training-free framework based on off-the-shelf retrievers and instruction-tuned LLMs that answers questions from an already-provided large context.
@@
To find relevant chunks with an imperfect forward-looking LLM, FB-RAG samples multiple outputs and assigns a high score to a chunk if it was used for any of them. This turns out to be powerful for improving RAG results over recent baselines across diverse tasks. We also find that gains can be achieved while reducing latency. On EN.QA dataset from Bench zhang-etal-2024-bench, one can combine a 70B parameter model for final response generation with an 8B model for forward-lookup and match the baseline performance with over % latency reduction. Further, one can get an % performance improvement with % latency reduction. Through our qualitative analysis, we find instances where even if all the sampled outputs from the smaller LLM incorrectly answer the input query, and often fail to follow our instructions properly, this still proves sufficient for the final, more powerful LLM to generate the correct response. We now summarize our contributions:
@@
We propose FB-RAG: a novel training-free framework for performing RAG with off-the-shelf instruction-tuned LLMs. FB-RAG employs a simple and effective look-ahead strategy to evaluate context chunks before selecting them for final response generation (Section 2).
@@
We comprehensively evaluate FB-RAG against recent training-free RAG and Long Context baselines on datasets from LongBench bai-etal-2024-longbench and Bench zhang-etal-2024-bench, finding that FB-RAG delivers consistent performance improvements. We further analyze key design choices in FB-RAG, such as the number of chunks retrieved and the number of samples used for forward lookup (Sections 3 and 4).
@@
We show that FB-RAG provides the flexibility to improve performance while also reducing the latency. We additionally perform qualitative analysis discussing the strengths and limitations of our approach, and provide insights for future progress in this area (Section 5).
@@ 2 Methodology
We focus on the task of question answering based on an already-provided context. Given an input query and a context , FB-RAG relies on an off-the-shelf retriever and instruction-tuned LLMs (without finetuning) to generate the output 111This general formulation encompasses several QA, summarization, and Multiple Choice Questions (MCQ) tasks - see Section 3 for the datasets considered in this work.. We assume that context is sufficient to answer the query , differentiating from some prior formulations that assume runtime access to web search engines yan2024corrective. At its core, FB-RAG relies on a look-ahead method to retrieve the most relevant context chunks from before performing the final response generation. We start by describing this method and later connect it to the overall three-stage process of FB-RAG.
@@ 2.1 Forward-Backward Retriever
Prior work has reported that LLMs often get confused by the irrelevant information present in the context xu2024recomp; asai2024self. The inverted U shape for the performance observed by yu2024defense as the context size increases demonstrates this in action. Hence, one obvious objective for the retrievers is to assign high importance scores to the most relevant chunks so that one can use a small context for generation and reduce irrelevant content. This is challenging for retrievers relying solely on the information in the input query, especially when the query is non-specific and complex li-etal-2024-retrieval. To address this gap, our key idea is to look forward at the potential answer to retrieve the relevant contexts. If we had access to the oracle generation model , we could compute in the following manner:
@@
Equation 5 presents the complete forward-looking component of our framework. In case of an extremely noisy model , the generated reasoning sequences and corresponding answers can be inaccurate and thus, can provide a misleading signal for our purpose of ranking the context chunks. Hence, merely relying on the outputs from such a noisy model can unfairly penalize the true relevant chunk . Motivated by this, we also incorporate a backward-looking component (as a form of a regularizer) that looks at the original input query to compute the importance scores:
where and denote the backward and forward components respectively, while and refer to their corresponding weights.
@@ 2.2 FB-RAG Overview
We evaluate two variants of our framework in this paper: 1) Ours-FB: Using both and as in Equation 2.1, and 2) Ours-F: Using and (ignoring the backward-component). As presented later in Sections 4 and 5, we find that Ours-F consistently outperforms Ours-FB across the board, indicating that one needs to only rely on the forward-looking component – at least for the choices for LLM considered in this work.
@@ 3 Experiment Design
Datasets: Following prior work li-etal-2024-retrieval, we focus on tasks that are a) in English, b) real, and c) query-based. This leads to datasets from LongBench bai-etal-2024-longbench: NarrativeQA kocisky-etal-2018-narrativeqa, Qasper dasigi-etal-2021-dataset, MultiFieldQA bai-etal-2024-longbench, HotpotQA yang-etal-2018-hotpotqa, 2WikiMultihopQA ho2020constructing, MuSiQue trivedi2022musique, and QMSum zhong-etal-2021-qmsum. We also pick two datasets from Bench zhang-etal-2024-bench, namely, En.QA and EN.MC. These datasets cover diverse domains, including Wikipedia articles, meetings, narratives, and research papers, involving single and multi-hop questions. The average context lengths range from a few thousand to k words. Refer to Appendix B for more details.
@@
Methods: Long Context (LC) refers to directly feeding the full context to the LLM without explicit retrieval. Vanilla denotes the typical RAG approach, which performs retrieval based on an off-the-shelf retriever before feeding the context to the LLM. We implemented two recent approaches evaluated on the considered datasets. In Order-Preserving (OP) RAG yu2024defense, the selected chunks from the retriever are first sorted in their original ordering before feeding them to the LLM. Self-Route li-etal-2024-retrieval is a look-ahead approach that relies on LLM’s ability to understand if the question is answerable from the retrieved context. It involves steps: 1) Retrieval: Based on an off-the-shelf retriever, 2) Generation: A modified generation based on the retrieved context where the LLM can choose to output ‘unanswerable’ if it finds that the retrieved context is insufficient to answer the question, and 3) Generation: Based on the full input context if the LLM outputs ‘unanswerable’ in the previous step.
For our approach, both Ours-FB and Ours-F variants use samples in Stage II obtained by combining top-p (p=) and top-k (k=) sampling. The final response generation for all methods uses Llama3.1-70B-Instruct llama31-2024. Self-Route uses the same model for both generation steps. For our approach, we use Llama3.1-8B-Instruct llama31-2024 for generating samples in Stage II. Refer to Appendix C for the prompts used, hardware details, and token limits. We evaluated retrievers: BM25 trotman2014improvements, M3Flag chen-etal-2024-m3, BGEFlag xiao2024c, and MPNet333https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1. We chose BM25 for our experiments due to its strong relative performance, simplicity, and versatility, making it suitable for our approach, which relies on LLM-generated outputs to retrieve relevant context chunks (see Appendix D.1 for a performance comparison). For chunking, we use a chunk size of words throughout.
@@ 4 Results
Self-Route li-etal-2024-retrieval | ||||||||
@@
OP yu2024defense | 44.34 | 23.89 | 49.31 | 54.8 | 61.11 | 59.06 | 37.94 | 24.26 |
Self-Route li-etal-2024-retrieval | 47.23 | 24.04 | 48.77 | 54.34 | 64.42 | 68.23 | 46.68 | 24.14 |
@@
OP yu2024defense | 48.03 | 26.62 | 50.71 | 56.78 | 66.28 | 64.8 | 45.91 | 25.11 |
Self-Route li-etal-2024-retrieval | 48.29 | 27.54 | 50.09 | 56.1 | 65.64 | 66.02 | 47.75 | 24.9 |
@@
OP li-etal-2024-retrieval | 48.75 | 29.85 | 51.35 | 55.6 | 65.53 | 65.5 | 48.85 | 24.59 |
Self-Route li-etal-2024-retrieval | 48.71 | 30.52 | 50.74 | 54.67 | 66.5 | 64.12 | 49.29 | 25.13 |
@@
Self-Route li-etal-2024-retrieval | ||
@@ 6 Related Work
Long Context (LC) LLMs: Context lengths have rapidly increased, with Gemini Pro team2024gemini and Meta Llama llama4-2025 using even million tokens. However, LLMs can get confused by irrelevant parts of the context, leading to known cases where RAG significantly outperforms LC xu2023retrieval; yu2024defense. In terms of latency, LC is expensive due to the quadratically increasing compute costs with input size. We follow the RAG paradigm by first retrieving the most relevant context chunks and then feeding them to an LLM with the input query for answer generation.
Retrieval Augmented Generation (RAG): RAG has emerged as a popular paradigm competing with LC, improving performance across diverse tasks with significantly lower compute costs fan2024survey. Traditional RAG is backward-looking – the context chunks are scored based on the input query using a combination of retrievers and rerankers, which further refine the selected context gao2023retrieval. Instead, FB-RAG uses forward-looking with samples generated from an LLM to select the relevant context chunks for the final answer generation. Unlike a typical reranker, Stage II of FB-RAG selects the chunks from the full context instead of (the output of Stage 1).
Numerous efforts augment RAG with trained filters yoran2023making, trained compressors xu2024recomp, and web search engines yan2024corrective to improve retrieval quality and generation. Self-RAG asai2024self trains an LLM using special reflection tokens to retrieve on demand. li2023web and jiang-etal-2023-active perform retrieval from the web based on the LLM’s look-ahead confidence scores. Speculative RAG wang2024speculative uses a smaller trained LLM to generate answer candidates, which are then verified by another LLM. LongRAG zhao2024longrag uses plug-n-play components to extract global information and factual details from context chunks which enhances the understanding from long contexts. Our setting differs in several ways: 1) We push the performance of instruction-tuned LLMs without any further training, 2) We assume no access to external web sources, and 3) We only use forward lookup from the light-weight LLM in a soft manner for selecting relevant context chunks from the entire context, with the final generation still being performed by a more powerful LLM. Two recent papers closest to our formulation are Self-Route li-etal-2024-retrieval and Order Preserving (OP) RAG yu2024defense, which we implemented ourselves and used as baselines in this work.

### v3

@@ Abstract
Traditional Retrieval-Augmented Generation (RAG) struggles with complex queries that lack strong signals to retrieve the most relevant context, forcing a trade-off between choosing a small context that misses key information and a large context that confuses the LLM. To address this, we propose Forward-Backward RAG (FB-RAG), a new training-free framework based on a simple yet powerful forward-looking strategy. FB-RAG employs a light-weight LLM to peek into potential future generations, using evidence from multiple sampled outputs to precisely identify the most relevant context for a final, more powerful generator. This improves performance without complex finetuning or Reinforcement Learning common in prior work. Across datasets from LongBench and Bench, FB-RAG consistently delivers strong results. Further, the performance gains can be achieved with reduced latency due to a shorter, more focused prompt for the powerful generator. On EN.QA dataset, FB-RAG matches the leading baseline with over % latency reduction or achieves an % performance improvement with a % latency reduction. Our analysis finds cases where even when the forward-looking LLM fails to generate correct answers, its attempts are sufficient to guide the final model to an accurate response, demonstrating how smaller LLMs can systematically improve the performance and efficiency of larger ones. Our code is available at: https://github.com/CapitalOne-Research/fb-rag.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) shows immense promise in reducing hallucinations and improving generation performance Fan et al. (2024); Gao et al. (2023). RAG achieves strong results on diverse Question Answering (QA) tasks Borgeaud et al. (2022); Guu et al. (2020); Asai et al. (2024), general language tasks He et al. (2021); Khandelwal et al. (2019), and across numerous downstream applications Liu et al. (2023); Wu et al. (2024).
@@
In this work, we focus on the task of answering queries based on an already-provided large context. Traditional RAG efforts for this setup involve two steps Zhao et al. (2024b): 1) Retrieve important chunks by computing similarities with the query (based on a sparse or dense retriever and/or a reranker), 2) Feed the retrieved chunks along with the query to an LLM, which generates the answer. We refer to these approaches as backward-looking – looking back at the input query to score the context chunks. Such methods have been widely adopted in both academia and industry. However, standard methods struggle with complex queries that lack sufficient information to retrieve relevant chunks (see example in Figure 1). This challenge is difficult to manage in RAG, where retrieving too little risks missing key information and retrieving too much risks adding irrelevant content that can confuse the LLMs Yu et al. (2024).
@@
To address this challenge, we design Forward-Backward RAG (FB-RAG) for studying the impact of an emerging yet underexplored idea – forward-looking or peeking into the LLM’s output generations to improve retrieval. FB-RAG generates the output in three stages: I) Recall-focused Retrieval, using an off-the-shelf retriever to extract a smaller, yet sufficiently large context, II) Precision-focused Retrieval, which either only relies on forward-looking by observing reasons and answers from a light-weight LLM to evaluate the context chunks (Ours-F) or relies on both forward and backward lookup (Ours-FB), and III) Generation, prompting a powerful LLM to get the final answer. Although prior work used related ideas to improve RAG with LLM-based feedback or confidence scores Zhao et al. (2024a); Sun et al. (2022); Wang et al. (2024); Yang et al. (2023); Jiang et al. (2023), these methods typically propose complex fine-tuning or Reinforcement Learning strategies, and often assume access to external web search engines or rely on LLM’s own memory which is not suitable for many domain-specific practical settings. Instead, FB-RAG is a simple and effective training-free framework based on off-the-shelf retrievers and instruction-tuned LLMs that answers questions from an already-provided large context.
@@
To find relevant chunks with an imperfect forward-looking LLM, FB-RAG samples multiple outputs and assigns a high score to a chunk if it was used for any of them. This turns out to be powerful for improving RAG results over recent baselines across diverse tasks. We also find that gains can be achieved while reducing latency. On EN.QA dataset from Bench Zhang et al. (2024), one can combine a 70B parameter model for final response generation with an 8B model for forward-lookup and match the baseline performance with over % latency reduction. Further, one can get an % performance improvement with % latency reduction. Through our qualitative analysis, we find instances where even if all the sampled outputs from the smaller LLM incorrectly answer the input query, and often fail to follow our instructions properly, this still proves sufficient for the final, more powerful LLM to generate the correct response. We now summarize our contributions:
@@
We propose FB-RAG: a training-free framework for performing RAG with off-the-shelf instruction-tuned LLMs. FB-RAG employs a simple and effective look-ahead strategy to evaluate context chunks before selecting them for final response generation (Section 2).
@@
We comprehensively evaluate FB-RAG against recent training-free RAG and Long Context baselines on datasets from LongBench Bai et al. (2024) and Bench Zhang et al. (2024), finding that FB-RAG delivers consistent performance improvements. We further analyze key design choices in FB-RAG, such as the number of chunks retrieved and the number of samples used for forward lookup (Sections 3 and 4).
@@
We show that FB-RAG provides the flexibility to improve performance while reducing latency. We additionally perform qualitative analysis discussing the strengths and limitations of our approach, and provide insights for future progress in this area (Section 5).
@@ 2 Methodology
We focus on the task of question answering based on an already-provided context. Given an input query and a context , FB-RAG relies on an off-the-shelf retriever and instruction-tuned LLMs (without finetuning) to generate the output 111This general formulation encompasses several QA, summarization, and Multiple Choice Questions (MCQ) tasks - see Section 3 for the datasets considered in this work.. We assume that context is sufficient to answer the query , differentiating from some prior formulations that assume runtime access to web search engines Yan et al. (2024). At its core, FB-RAG relies on a look-ahead method to retrieve the most relevant context chunks from before performing the final response generation. We start by describing this method and later connect it to the overall three-stage process of FB-RAG.
@@ 2.1 Forward-Backward Retriever
Prior work has reported that LLMs often get confused by the irrelevant information present in the context Xu et al. (2024); Asai et al. (2024). The inverted U shape for the performance observed by Yu et al. (2024) as the context size increases demonstrates this in action. Hence, one obvious objective for the retrievers is to assign high importance scores to the most relevant chunks so that one can use a small context for generation and reduce irrelevant content. This is challenging for retrievers relying solely on the information in the input query, especially when the query is non-specific and complex Li et al. (2024). To address this gap, our key idea is to look forward at the potential answer to retrieve the relevant contexts. If we had access to the oracle generation model , we could compute in the following manner:
@@
Equation 5 utilizes the complete forward-looking component of our framework (referred as below). However, our formulation of is still incomplete. In case of an extremely noisy model , the generated reasoning sequences and corresponding answers can be inaccurate and thus, can provide a misleading signal for our purpose of ranking the context chunks. Hence, merely relying on the outputs from such a noisy model can unfairly penalize the true relevant chunk . Motivated by this, we also incorporate a backward-looking component (as a form of a regularizer) that looks at the original input query to compute the importance scores:
where and denote the backward and forward components respectively, while and refer to their corresponding weights. Equation 2.1 completes our goal for this Section.
@@ 2.2 FB-RAG Overview
We evaluate two variants of FB-RAG in this paper: 1) Ours-FB: Using both and as in Equation 2.1, and 2) Ours-F: Using and (ignoring the backward-component and resorting to the formulation in Equation 5 instead). As presented later in Sections 4 and 5, we find that Ours-F consistently outperforms Ours-FB across the board, indicating that one needs to only rely on the forward-looking component – at least for the choices for LLM considered in this work.
@@ 3 Experiment Design
Datasets: Following prior work Li et al. (2024), we focus on tasks that are a) in English, b) real, and c) query-based. This leads to datasets from LongBench Bai et al. (2024): NarrativeQA Kočiský et al. (2018), Qasper Dasigi et al. (2021), MultiFieldQA Bai et al. (2024), HotpotQA Yang et al. (2018), 2WikiMultihopQA Ho et al. (2020), MuSiQue Trivedi et al. (2022), and QMSum Zhong et al. (2021). We also pick two datasets from Bench Zhang et al. (2024), namely, En.QA and EN.MC. These datasets cover diverse domains, including Wikipedia articles, meetings, narratives, and research papers, involving single and multi-hop questions. The average context lengths range from a few thousand to k words. Refer to Appendix B for more details.
@@
Methods: Long Context (LC) refers to directly feeding the full context to the LLM without explicit retrieval. Vanilla denotes the typical RAG approach, which performs retrieval based on an off-the-shelf retriever before feeding the context to the LLM. We implemented two recent approaches evaluated on the considered datasets. In Order-Preserving (OP) RAG Yu et al. (2024), the selected chunks from the retriever are first sorted in their original ordering before feeding them to the LLM. Self-Route Li et al. (2024) is a look-ahead approach that relies on LLM’s ability to understand if the question is answerable from the retrieved context. It involves steps: 1) Retrieval: Based on an off-the-shelf retriever, 2) Generation: A modified generation based on the retrieved context where the LLM can choose to output ‘unanswerable’ if it finds that the retrieved context is insufficient to answer the question, and 3) Generation: Based on the full input context if the LLM outputs ‘unanswerable’ in the previous step.
For our approach, both Ours-FB and Ours-F variants use samples in Stage II obtained by combining top-p (p=) and top-k (k=) sampling. The final response generation for all methods uses Llama3.1-70B-Instruct Meta (2024). Self-Route uses the same model for both generation steps. For our approach, we use Llama3.1-8B-Instruct Meta (2024) for generating samples in Stage II. Refer to Appendix C for the prompts used, hardware details, and token limits. We evaluated retrievers: BM25 Trotman et al. (2014), M3Flag Chen et al. (2024), BGEFlag Xiao et al. (2024), and MPNet333https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1. We chose BM25 for our experiments due to its strong relative performance, simplicity, and versatility, making it suitable for our approach, which relies on LLM-generated outputs to retrieve relevant context chunks (see Appendix D.1 for a performance comparison). For chunking, we use a chunk size of words throughout.
@@ 4 Results
| Self-Route Li et al. (2024) | ||||||||
@@
| OP Yu et al. (2024) | 44.34 | 23.89 | 49.31 | 54.8 | 61.11 | 59.06 | 37.94 | 24.26 |
| Self-Route Li et al. (2024) | 47.23 | 24.04 | 48.77 | 54.34 | 64.42 | 68.23 | 46.68 | 24.14 |
@@
| OP Yu et al. (2024) | 48.03 | 26.62 | 50.71 | 56.78 | 66.28 | 64.8 | 45.91 | 25.11 |
| Self-Route Li et al. (2024) | 48.29 | 27.54 | 50.09 | 56.1 | 65.64 | 66.02 | 47.75 | 24.9 |
@@
| OP Li et al. (2024) | 48.75 | 29.85 | 51.35 | 55.6 | 65.53 | 65.5 | 48.85 | 24.59 |
| Self-Route Li et al. (2024) | 48.71 | 30.52 | 50.74 | 54.67 | 66.5 | 64.12 | 49.29 | 25.13 |
@@
| Self-Route Li et al. (2024) | ||
@@ 5 Discussion
Combining FB-RAG with In-Context Learning: Wei et al. (2024) proposed to improve the generation step in RAG with In-Context Learning (ICL) examples. This approach is complementary to FB-RAG, which improves the retrieval of relevant chunks. Hence, we studied how combining these two methods impacts performance. We consider two variants for the retrieval step: 1) Vanilla and 2) Ours-F). We consider three augmentations: Reason-then-Answer, which includes reasoning in the final generation by Llama3.1-70B-Instruct, along with Few-Shot Demo. w/ Instruction and INSTRUCTRAG-ICL methods by Wei et al. (2024) which combine reasoning with ICL. We implemented these methods ourselves and combined them with one of the two retrieval choices, for a total of new evaluations.
We present the results in Appendix Table 9. On average, our original Ours-F model from Table 1 remains the top performer. The ICL augmentations provide a clear benefit on one dataset (2Wiki). Crucially, the results for the ICL-augmented Vanilla setting are consistently lower than for the ICL-augmented Ours-F. This demonstrates that while generation techniques like ICL can help, they cannot fully compensate for poor retrieval. Ours-F provides a better, more precise context for these generation methods to work with.
Our qualitative analysis reveals trade-offs. Adding reasoning (especially Reason-then-Answer) often hurts instruction-following, causing the model to generate long answers instead of the concise phrases requested in the prompt. While INSTRUCTRAG-ICL fixes some errors made by the base Ours-F model, its own faulty reasoning sometimes introduces new errors. One potential benefit we observed was in handling out-of-context queries: the INSTRUCTRAG-ICL model often correctly concluded that information was missing or used its internal knowledge, whereas the base Ours-F model was more likely to make an incorrect guess. While this did not improve scores after the augmentations, it suggests a potential direction for improving system reliability. Effectively integrating these ICL methods remains an area for future work.
@@ 6 Related Work
Long Context (LC) LLMs: Context lengths have rapidly increased, with Gemini Pro Team et al. (2024) and Meta Llama Meta (2025) using even million tokens. However, LLMs can get confused by irrelevant parts of the context, leading to known cases where RAG significantly outperforms LC Xu et al. (2023); Yu et al. (2024). In terms of latency, LC is expensive due to the quadratically increasing compute costs with input size. We follow the RAG paradigm by first retrieving the most relevant context chunks and then feeding them to an LLM with the input query for answer generation.
Retrieval Augmented Generation (RAG): RAG has emerged as a popular paradigm competing with LC, improving performance across diverse tasks with significantly lower compute costs Fan et al. (2024). Traditional RAG is backward-looking – the context chunks are scored based on the input query using a combination of retrievers and rerankers, which further refine the selected context Gao et al. (2023). Instead, FB-RAG uses forward-looking with samples generated from an LLM to select the relevant context chunks for the final answer generation. Unlike a typical reranker, Stage II of FB-RAG selects the chunks from the full context instead of (the output of Stage 1).
Numerous efforts augment RAG with trained filters Yoran et al. (2023), trained compressors Xu et al. (2024), and web search engines Yan et al. (2024) to improve retrieval quality and generation. Self-RAG Asai et al. (2024) trains an LLM using special reflection tokens to retrieve on demand. Li et al. (2023) and Jiang et al. (2023) perform retrieval from the web based on the LLM’s look-ahead confidence scores. Speculative RAG Wang et al. (2024) uses a smaller trained LLM to generate answer candidates, which are then verified by another LLM. LongRAG Zhao et al. (2024a) uses plug-n-play components to extract global information and factual details from context chunks which enhances the understanding from long contexts. Our setting differs in several ways: 1) We push the performance of instruction-tuned LLMs without any further training, 2) We assume no access to external web sources, and 3) We only use forward lookup from the light-weight LLM in a soft manner for selecting relevant context chunks from the entire context, with the final generation still being performed by a more powerful LLM. Two recent papers closest to our formulation are Self-Route Li et al. (2024) and Order Preserving (OP) RAG Yu et al. (2024), which we implemented ourselves and used as baselines in this work.